# Sequential Monte Carlo with Gaussian Mixture Approximation for Infinite-Dimensional Statistical Inverse Problems

### v3

@@ Abstract
By formulating the inverse problem of partial differential equations (PDEs) as a statistical inference problem, the Bayesian approach provides a general framework for quantifying uncertainties. In the inverse problem of PDEs, parameters are defined on an infinite-dimensional function space, and the PDEs induce a computationally intensive likelihood function. Additionally, sparse data tends to lead to a multi-modal posterior. These features make it difficult to apply existing sequential Monte Carlo (SMC) algorithms. To overcome these difficulties, we propose new conditions for the likelihood functions, construct a Gaussian mixture based preconditioned Crank-Nicolson transition kernel, and demonstrate the universal approximation property of the infinite-dimensional Gaussian mixture probability measure. By combining these three novel tools, we propose a new SMC algorithm, named SMC-GM. For this new algorithm, we obtain a convergence theorem that allows Gaussian priors, illustrating that the sequential particle filter actually reproduces the true posterior distribution. Furthermore, the proposed new algorithm is rigorously defined on the infinite-dimensional function space, naturally exhibiting the discretization-invariant property. Numerical experiments demonstrate that the new approach has a strong ability to probe the multi-modality of the posterior, significantly reduces the computational burden, and numerically exhibits the discretization-invariant property (important for large-scale problems).
@@ 1 Introduction
The MCMC method, serving as a vital component within the Bayesian framework, provides the means to navigate and sample from these intricate posterior distributions. In the infinite-dimensional setting, the preconditioned Crank-Nicolson (pCN) and pCN Langevin transition kernel can be derived from the discretization of the Langevin equation Andrew_review , BESKOS2008 , Hairer2007 . However, the Bayesian inference for computationally intensive models can be limited by the expense of MCMC sampling. In theory, MCMC methods can sample from the posterior distribution, but in practice, the computational cost of MCMC in PDE inverse problems can be very high. Additionally, the burn-in process and the difficulty of parallelizing traditional MCMC YeGuifeng2024 further contribute to the computational expense.
@@
    Use a surrogate or approximate likelihood to avoid evaluating the expensive likelihood when possible Bon2021 .
@@
    Combine other useful tools with MCMC, such as variational inference Naesseth2018 .
@@
The first strategy has an extremely wide range of applications. The choices of transition kernels in SMC methods are highly flexible. For example, the basic random walk and its infinite-dimensional improvement, pCN, are used in SMC as described in Kantas2014 and Beskos2015SC . The dimension-independent likelihood-informed transition kernel is utilized in Cui2024 . Another SMC method that leverages the gradient information of the likelihood function is presented in Rosato2024 . Finally, a series of studies Gunawan2020 , Buchholz2021 , Martin2023 employs Hamiltonian Monte Carlo as the transition kernel to enhance sampling efficiency.
In this paper, we focus on improving SMC in infinite-dimensional spaces. To further reduce computational costs, we consider improvements for the MCMC transition steps. SMC methods are particularly attractive for a diverse range of inference challenges due to their ability to effectively manage multi-modal distributions. We aim to preserve these desirable approximation properties while performing the approximation. For handling multi-modal distributions, the Gaussian mixture is an appropriate choice. In fact, it has been proven that a Gaussian mixture can approximate any distribution in the finite-dimensional case Nguyen2020approximation . The Gaussian mixture approximation, employed in particle filters Raihan2018 , is a commonly used method for approximation. Gaussian mixture approximations have been applied in SMC filtering. In YunSehyun2019 , the authors constructed the prior samples into a Dirac measure, regarded it as a Gaussian mixture distribution with zero variance, and then linearized the model to update the parameters of the Gaussian mixture, ultimately obtaining an approximation of the posterior.
Unlike these works, we will combine the first two strategies by incorporating a likelihood-informed Gaussian mixture into the pCN proposal, resulting in the pCN-GM method. Furthermore, we can utilize a Gaussian mixture sampler to supplant the mutation step, dubbing it the sequential Monte Carlo algorithm with Gaussian mixture approximation (SMC-GM). The main difficulties are as follows: First, existing SMC convergence theorems require the potential function to be bounded Beskos2015SC , yet commonly used Gaussian priors and many forward problems, such as Darcy flow, cannot apply this theorem. Second, traditional MCMC simulators using Metropolis-Hastings do not necessarily behave well when dealing with sampling from multi-modal distributions Daviet2018 . Additionally, in the infinite-dimensional setting, there are singularity issues between measures, such as changes of the mean of a Gaussian measure, which almost surely lead to singularities Bogachev1998 . Third, while mixture Gaussian densities can approximate density functions well in finite-dimensional settings, can infinite-dimensional Gaussian mixture measures also approximate the posterior measure, and if so, in what sense of distance?
@@
    We derive a generalized pCN algorithm based on Gaussian mixtures (pCN-GM) from the Crank-Nicolson discretization of the Langevin system that describes multiple particles with various perturbations. This prevents the occurrence of singularities, i.e., the pCN-GM method is well-defined under the infinite-dimensional setting. This algorithm can be viewed as an extension of both pCN and independent sampling with Gaussian mixtures fengzhe2018 .
@@
    Based on the fact that finite-dimensional Gaussian mixture density functions can approximate any density function, along with the continuity of the likelihood function and the properties of infinite-dimensional Gaussian measures, we prove the denseness of Gaussian mixture measures in infinite-dimensional spaces. This provides a solid foundation for the SMC-GM algorithm.
This paper is organized as follows. In Section 2 we construct the SMC-GM method. Subsection 2.1 provides a brief introduction to the SMC method. In Subsection 2.2, we prove the convergence theorem for SMC under weak conditions. In Subsection 2.3 we propose a new Markov transition kernel and prove that it is well-defined in the infinite-dimensional space. In Subsection 2.4 we present the universal approximation property of Gaussian mixture measures in the infinite-dimensional setting, and give the approximated SMC-GM method. Finally, Section 3 applies the developed SMC-GM method to three typical inverse problems. All proofs are provided in section 5.4 in the supplementary material. All the programming codes are accessible at https://github.com/jjx323/IPBayesML-SMC-GM.
@@ 2 SMC with Gaussian mixture approximation
In this section, we adopt the SMC methods defined in infinite-dimensional function spaces Beskos2015SC and implement three critical improvements. First, we provide a convergence theorem for SMC under weaker conditions. Second, we combine the Gaussian mixture with pCN to introduce a new transition kernel, known as the pCN-GM algorithm, which, when utilized as the transition kernel in SMC, forms the SMC-pCN-GM method. Lastly, we present an approximation to SMC-pCN-GM, the SMC-GM algorithm, and subsequent numerical experiments will demonstrate that this approximation significantly enhances efficiency.
@@ 2.1 Algorithm overview
Now we provide a brief introduction to the SMC method, for more details, see Moral2006 , Beskos2015SC . Lastly, we propose the SMC-GM algorithm, an efficient approximation of the SMC method.
Recall that $\mu_{0}$ is the prior measure on a separable Hilbert space $\mathcal{H}$, equipped with the Borel $\sigma$-algebra $\mathcal{B}(\mathcal{H})$. We aim to sample from the posterior measure $\mu^{\mathrm{d}}$ defined in (2.4). The SMC methods share the general structure of particle filters, which uses a $N$-particle Dirac measure
@@
where $\{h_{j}\}_{j=1}^{J}$ is a positive sequence satisfying $\sum_{j=1}^{J}h_{j}=1$. The method for determining $h_{i}$ is described in section 5.3 in the supplementary materials. Now we have a measure sequences $\{\mu_{j}\}_{j=0}^{J}$, and the importance sampling can be used to sample from $\mu_{j+1}$ based on samples from $\mu_{j}$.
@@
For readers’ convenience, we detail the aforementioned transition kernels (pCN, pCN-GM, GM, RW) in section 5.1 in the supplementary material. To adapt an SMC-XX method, simply replace the mutation step in Algorithm 1 with the XX method.
@@ 2.2 Approximation theory of SMC
(a). They are lower bounded, i.e., there exists $\kappa_{1}>0$ such that for all $j\in\{1,2,\cdots,$J}J\}, we have
@@
(b). They are finite almost everywhere, i.e., for all $j\in\{1,2,\cdots,J\}$ we have
@@
(c). The sum $\Phi=\sum_{j=1}^{J}\Phi_{j}$ satisfies the following integral property: there exists $\kappa_{2}>0$ such that
$Z_{J}\mathrel{\mathop{\ordinarycolon}}=\int_{\mathcal{H}}\exp\left(-\Phi(u)\right)\mu_{0}(du)>\kappa_{2}>0.$
@@
The final operator is the sampling operator, because we use a weighted sum of Dirac measure to approximate the posterior:
@@
Using these three notations, the measure of the next layer in SMC can be expressed in terms of the measure of the previous layer as follows:
$\mu_{j+1}^{N}=L_{j+1}S^{N}Q_{j}\mu_{j}^{N},\quad 0\leq j\leq J-1,\quad\mu_{0}^{N}=\mu_{0}.$ (2.11)
@@
which is the total variation distance. For more details, see section 5.4.1 in the supplementary material.
We can estimate the error induced by the above three operators $L_{j},P_{j}$, and $S^{N}$ defined in (2.8),(2.7), and (2.10):
@@
(a). The sampling operator $S^{N}$ satisfies
@@
(b). The Markov kernel $P_{j}$ satisfies
@@
(c). Under Assumptions 2.1, let $\nu_{j}=P_{j}\mu_{j},$ then we have
$d(L_{j}\nu_{j-1},L_{j}\mu)\leq 4\kappa_{1}^{-1}\kappa_{2}^{-1}d(\nu_{j-1},\mu).$ (2.16)
@@
Proof of Lemma 2.1.(a),(b), see Dashti2017 , and proof of Lemma 2.1.(c), see section 5.4.2 in the supplementary material.
@@
$d(\mu_{J}^{N},\mu_{J})$ $\leq\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\left(\frac{4}{\kappa_{1}\kappa_{2}}\right)^{j}.$
Theorem 2.2 ensures the rationality of SMC for posterior sampling derived from a Gaussian prior. The proof is provided in section 5.4.3 in the supplementary material. In the next subsection, we will refine the pCN algorithm, which is also designed for posteriors derived from a Gaussian prior. Without Theorem 2.2, using pCN as a transition kernel would lack theoretical justification.
@@
The key components of this kernel are the proposal $Q(u,dv)$ and the acceptance rate function $a(u,v)$. To sample from the posterior, we must select $Q(u,dv)$ and $a(u,v)$ such that $P(u,dv)$ defined by (2.17) preserves $\mu^{\mathrm{d}}$ as follows:
$\int_{u\in\mathcal{H}}P(u,dv)\mu^{\mathrm{d}}(du)=\mu^{\mathrm{d}}(dv).$ (2.18)
@@
$a_{\text{\tiny MH}}=\min\left\{1,\frac{\mu^{\mathrm{d}}(dv)Q(v,du)}{\mu^{\mathrm{d}}(du)Q(u,dv)}\right\},$ (2.19)
so that $P(u,dv)$ preserves $\mu^{\mathrm{d}}$, provided that $\mu^{\mathrm{d}}(du)Q(u,dv)$ and $\mu^{\mathrm{d}}(dv)Q(v,du)$ are equivalent as measures on $(\mathcal{H}\times\mathcal{H},B(\mathcal{H})\times B(\mathcal{H}))$. We will refer to a Metropolis-Hastings algorithm as well-defined if this condition of absolute continuity is satisfied CUI2016 . Hence, we only need to select the proposal kernel $Q(u,dv)$ and demonstrate the equivalence between $\mu^{\mathrm{d}}(du)Q(u,dv)$ and $\mu^{\mathrm{d}}(dv)Q(v,du)$.
@@
where $\mathcal{N}_{0,\mathcal{C}}$ denotes the Gaussian measure $\mathcal{N}(0,\mathcal{C})$. Note that (2.20) is a drifted version pCNwithmean of the pCN proposal. The special case without drift (i.e., when $m=0$) is more frequently discussed Andrew_review , Beskos2015SC . This kernel allows for deriving a simple acceptance rate function
$a_{\text{pCN}}(u,v)=\min\left\{1,\exp(\Phi(u)-\Phi(v))\right\}.$ (2.21)
The pCN method will serve as the transition kernel for SMC in numerical experiments, constituting an accurate posterior sampling algorithm for comparison with the SMC-GM algorithm. On the other hand, to prove the convergence of our algorithm (in Section 2.3), we generalize the pCN method to the case with a Gaussian mixture proposal, and prove that it is also well-defined in an infinite-dimensional space.
@@
This proposal has been carefully constructed, and it will yield a well-defined Metropolis-Hastings algorithm at the end of this subsection. We can obtain this proposal from the Crank-Nicolson discretization of a Langevin system, see section 5.2 in the supplementary material. This proposal degenerates to the pCN proposal with a non-zero mean prior pCNwithmean when $M=1$, and to the proposal in the Gaussian mixture independence sampler fengzhe2018 when $\beta=1$. We will estimate the parameters of a Gaussian mixture from samples by the method from fengzhe2018 , which is established in function space. And the resulting Gaussian mixture is also equivalent to the prior measure $\mu_{0}$.
@@
$a(u,v)=\min\left\{1,\frac{\mu^{\mathrm{d}}(dv)Q(v,du)}{\mu^{\mathrm{d}}(du)Q(u,dv)}\right\}.$ (2.25)
Here the accept rate function $a(u,v)$ is well-defined in infinite-dimensional Hilbert space if and only if the Radon-Nikodym derivative exists, that is, $\mu^{\mathrm{d}}(dv)Q(v,du)$ is absolutely continuous with respect to $\mu^{\mathrm{d}}(du)Q(u,dv)$, as measures on $\mathcal{H}\times\mathcal{H}$. We will first illustrate the well-definiteness of the measure $\mu_{0}(dv)Q(v,du)$ and then show their equivalence.
@@
$\mu_{0}(du)Q(u,dv)=$ $\sum_{j=1}^{M}w_{j}\mu_{j}(du,dv),\quad\mu_{j}=\mathcal{N}\left(m_{j},\mathcal{V}_{j}\right),$ (2.26)
$\mu_{0}(dv)Q(v,du)=$ $\sum_{j=1}^{M}w_{j}\mu_{j}^{\prime}(du,dv),\quad\mu_{j}^{\prime}=\mathcal{N}\left(m_{j}^{\prime},\mathcal{V}_{j}^{\prime}\right).$ (2.27)
@@
The proof of Theorem 2.3 is given in section 5.4.4 in the supplementary material. Furthermore we can prove the equivalence between $\mu_{0}(dv)Q(v,du)$ and $\mu_{0}(du)Q(u,dv).$
@@
Under all of the assumptions of Theorem 2.3, the operator $\mathcal{V}_{j}^{-1/2}\mathcal{V}_{j}^{\prime}\mathcal{V}_{j}^{-1/2}$-$\mathcal{I}$ is a Hilbert-Schmidt operator, and $\mu_{0}(du)Q(u,dv)$ is equivalent to $\mu_{0}(dv)Q(v,du)$. In fact, all of their components are equivalent. Furthermore, $\mu^{\mathrm{d}}(du)Q(u,dv)$ and $\mu^{\mathrm{d}}(dv)Q(v,du)$ are equivalent.
@@
The equivalence between the two measures in this theorem is crucial(see the proof in section 5.4.5 in the supplementary material), directly affecting whether the algorithm possesses dimension independence. A simple example is the comparison between the pCN and the random walk transition kernel:
@@
where just a slight difference in coefficients can lead to significant changes in the algorithm’s acceptance rate as the discrete dimension increasesAndrew_review . Moreover, if we require a proposal of the form given in Equation (2.24) without the assumption $\gamma=\sqrt{1-\beta^{2}}$ to yield a well-defined acceptance rate function, then the only option is $\beta^{2}+\gamma^{2}=1$. Discarding the meaningless negative solutions, the generalization we provide is unique. The proof of uniqueness is given in section 5.4.6 the supplementary material.

### v4

@@ Abstract
By formulating the inverse problem of partial differential equations (PDEs) as a statistical inference problem, the Bayesian approach provides a general framework for quantifying uncertainties. In the inverse problem of PDEs, parameters are defined on an infinite-dimensional function space, and the PDEs induce a computationally intensive likelihood function. Additionally, sparse data tends to lead to a multi-modal posterior. These features make it difficult to apply existing sequential Monte Carlo (SMC) algorithms. To overcome these difficulties, we propose new conditions for the likelihood functions, construct a Gaussian mixture based preconditioned Crank-Nicolson transition kernel, and demonstrate the universal approximation property of the infinite-dimensional Gaussian mixture probability measure. By combining these three novel tools, we propose a new SMC algorithm with Gaussian mixture approximation, together with an easy-to-use reduced version. For this new algorithm, we obtain a convergence theorem that allows Gaussian priors, illustrating that the sequential particle filter actually reproduces the true posterior distribution. Furthermore, the proposed new algorithm is rigorously defined on the infinite-dimensional function space, naturally exhibiting the discretization-invariant property. Numerical experiments demonstrate that the reduced version has a strong ability to probe the multi-modality of the posterior, significantly reduces the computational burden, and numerically exhibits the discretization-invariant property (important for large-scale problems).
@@ 1 Introduction
The MCMC method, serving as a vital component within the Bayesian framework, provides the means to navigate and sample from these intricate posterior distributions. In the infinite-dimensional setting, the preconditioned Crank-Nicolson (pCN) and pCN Langevin transition kernel can be derived from the discretization of the Langevin equation Andrew_review , Beskos2008 , Hairer2007 . However, the Bayesian inference for computationally intensive models can be limited by the expense of MCMC sampling. In theory, MCMC methods can sample from the posterior distribution, but in practice, the computational cost of MCMC in PDE inverse problems can be very high. Additionally, the burn-in process and the difficulty of parallelizing traditional MCMC Ye2024 further contribute to the computational expense.
@@
    Use a surrogate or approximate likelihood to avoid the expensive computation cost if possible Bon2021 .
@@
    Combine other useful tools with MCMC, such as variational inference Naesseth2017 .
@@
The first strategy has an extremely wide range of applications. The choices of transition kernels in SMC methods are highly flexible. For example, the basic random walk and its infinite-dimensional improvement, pCN, are used in SMC as described in Kantas2014 and Beskos2015SC . The dimension-independent likelihood-informed transition kernel is utilized in Cui2024 . Another SMC method that leverages the gradient information of the likelihood function is presented in Rosato2024 . Finally, a series of studies Gunawan2020 , Buchholz2021 , Burda2023 employs Hamiltonian Monte Carlo as the transition kernel to enhance sampling efficiency.
In this paper, we focus on improving SMC in infinite-dimensional spaces. To further reduce computational costs, we consider improvements for the MCMC transition steps. SMC methods are particularly attractive for a diverse range of inference challenges due to their ability to effectively manage multi-modal distributions. We aim to preserve these desirable approximation properties while performing the approximation. For handling multi-modal distributions, the Gaussian mixture is an appropriate choice. In fact, it has been proven that a Gaussian mixture can approximate any distribution in the finite-dimensional case Nguyen2020 . The Gaussian mixture approximation, employed in particle filters Raihan2018 , is a commonly used method for approximation. Gaussian mixture approximations have been applied in SMC filtering. In Sehyun2019 , the authors constructed the prior samples into a Dirac measure, regarded it as a Gaussian mixture distribution with zero variance, and then linearized the model to update the parameters of the Gaussian mixture, ultimately obtaining an approximation of the posterior.
Unlike these works, we will combine the first two strategies by incorporating a likelihood-informed Gaussian mixture into the pCN proposal, resulting in the pCN-GM method. Employing pCN-GM as the mutation step in SMC leads to the SMC-pCN-GM algorithm. Furthermore, by replacing the mutation step with a Gaussian mixture sampler, we obtain a reduced version, termed SMC-GM.
The main difficulties are as follows: First, existing SMC convergence theorems require the potential function to be bounded Beskos2015SC , yet commonly used Gaussian priors and many forward problems, such as Darcy flow, cannot apply this theorem. Second, traditional MCMC simulators using Metropolis-Hastings do not necessarily behave well when dealing with sampling from multi-modal distributions Daviet2018 . Additionally, in the infinite-dimensional setting, there are singularity issues between measures, such as changes of the mean of a Gaussian measure, which almost surely lead to singularities Bogachev1998 . Third, while mixture Gaussian densities can approximate density functions well in finite-dimensional settings, can infinite-dimensional Gaussian mixture measures also approximate the posterior measure, and if so, in what sense of distance?
@@
    We derive a generalized pCN algorithm based on Gaussian mixtures (pCN-GM) from the Crank-Nicolson discretization of the Langevin system that describes multiple particles with various perturbations. This prevents the occurrence of singularities, i.e., the pCN-GM method is well-defined under the infinite-dimensional setting. This algorithm can be viewed as an extension of both pCN and independent sampling with Gaussian mixtures Feng2018 .
@@
    Using pCN-GM in the mutation step of SMC yields the SMC-pCN-GM method. Based on the fact that finite-dimensional Gaussian mixture density functions can approximate any density function, along with the continuity of the likelihood function and the properties of infinite-dimensional Gaussian measures, we prove the denseness of Gaussian mixture measures in infinite-dimensional spaces. This motivates a reduced version of the SMC-pCN-GM method.
This paper is organized as follows. In Section 2 we propose the SMC-pCN-GM method, together with its reduced version, SMC-GM. Both methods are SMC algorithms with Gaussian mixture approximation. Subsection 2.1 provides a brief introduction to the SMC method. In Subsection 2.2, we prove the convergence theorem for SMC under weak conditions. In Subsection 2.3 we propose the pCN-GM kernel, a new Markov transition kernel, and prove that it is well-defined in the infinite-dimensional space. In Subsection 2.4 we present the universal approximation property of Gaussian mixture measures in the infinite-dimensional setting, and give the approximated SMC-GM method. Finally, Section 3 applies the developed SMC-GM method to three typical inverse problems. All proofs are provided in 5.4. All the programming codes are accessible at https://github.com/jjx323/IPBayesML-SMC-GM.
@@ 2 SMC with Gaussian mixture approximation
In this section, we adopt the SMC methods defined in infinite-dimensional function spaces Beskos2015SC and implement three critical improvements. First, we provide a convergence theorem for SMC under weaker conditions. Second, we combine the Gaussian mixture with pCN to introduce a new transition kernel, known as the pCN-GM algorithm, which, when utilized as the transition kernel in SMC, forms the SMC-pCN-GM method. Lastly, we present an approximation to SMC-pCN-GM as a reduced version, the SMC-GM algorithm.
@@ 2.1 Algorithm overview
Now we provide a brief introduction to the SMC method, for more details, see Moral2006 , Beskos2015SC . Recall that $\mu_{0}$ is the prior measure on a separable Hilbert space $\mathcal{H}$, equipped with the Borel $\sigma$-algebra $\mathcal{B}(\mathcal{H})$. We aim to sample from the posterior measure $\mu^{\mathrm{d}}$ defined in (2.4). The SMC methods share the general structure of particle filters, which uses a $N$-particle Dirac measure
@@
where $\{h_{j}\}_{j=1}^{J}$ is a positive sequence satisfying $\sum_{j=1}^{J}h_{j}=1$. The method for determining $h_{i}$ is described in 5.3. Now we have a measure sequences $\{\mu_{j}\}_{j=0}^{J}$, and the importance sampling can be used to sample from $\mu_{j+1}$ based on samples from $\mu_{j}$.
@@
For readers’ convenience, we detail the aforementioned transition kernels (pCN, pCN-GM, GM, RW) in 5.1. To adapt an SMC-XX method, simply replace the mutation step in Algorithm 1 with the XX method. The algorithms proposed in this paper are SMC methods with Gaussian mixture approximations, including SMC-pCN-GM and its reduced version SMC-GM.
@@ 2.2 Approximation theory of SMC
(a) They are lower bounded, i.e., there exists $\kappa_{1}>0$ such that for all $j\in\{1,2,\cdots,$J}J\}, we have
@@
(b) They are finite almost everywhere, i.e., for all $j\in\{1,2,\cdots,J\}$ we have
@@
(c) There exists $\kappa_{2,j}>0$ such that
$Z_{j}\mathrel{\mathop{\ordinarycolon}}=\int_{\mathcal{H}}\exp\left(-\Phi_{j}(u)\right)\mu_{j-1}(du)>\kappa_{2,j-1},\quad 1\leq j\leq J.$
@@
The third operator is the sampling operator, because we use a weighted sum of Dirac measure to approximate the posterior:
@@
The final operator is the resampling operator:
$R\!\left(\sum_{n=1}^{N}w(v^{(n)})\delta(v-v^{(n)})\right)$ $=\frac{1}{N}\sum_{n=1}^{N}\delta(v-\tilde{v}^{(n)}),\quad\tilde{v}^{(n)}\sim\sum_{n=1}^{N}w(v^{(n)})\delta(v-v^{(n)}).$
Using these notations, the measure of the next layer in SMC can be expressed in terms of the measure of the previous layer as follows:
$\mu_{j+1}^{N}=RL_{j+1}S^{N}Q_{j}\mu_{j}^{N},\quad 0\leq j\leq J-1,\quad\mu_{0}^{N}=\mu_{0}.$ (2.11)
@@
which is the total variation distance. For more details, see 5.4.1.
Let $\mathcal{P}(\mathcal{H})$ denote the set of all probability measures on $\mathcal{H}$. We can estimate the error induced by the above three operators $L_{j},P_{j}$, and $S^{N}$ defined in (2.8),(2.7), and (2.10):
@@
(a) The sampling operator $S^{N}$ satisfies
@@
(b) For any $\nu,\nu^{\prime}\in\mathcal{P}(\mathcal{H})$, the Markov kernel $P_{j}$ satisfies
@@
(c) Under Assumptions 2.1, for any $\mu\in\mathcal{P}(\mathcal{H})$, we have
$d(L_{j}\mu_{j-1},L_{j}\mu)\leq 2\kappa_{1}^{-1}\kappa_{2,j-1}^{-1}d(\nu_{j-1},\mu).$ (2.16)
(d) Let $\mu^{N}=\sum_{n=1}^{N}w(v^{(n)})\delta_{v^{(n)}}$, then we have
$d(R\mu^{N},\mu^{N})\leq\frac{1}{\sqrt{N}}.$
@@
Proof of Lemma 2.1.(a),(b), see Dashti2017 , and proof of Lemma 2.1.(c),(d), see 5.4.2, 5.4.3.
@@
$d(\mu_{J}^{N},\mu_{J})\leq\frac{1}{\sqrt{N}}\Biggl[1+2\sum_{j=1}^{J-1}\frac{2^{j}}{\kappa_{1}^{j}}\prod_{k=1}^{j}\frac{1}{\kappa_{2,J-k}}+\prod_{j=0}^{J-1}\frac{2}{\kappa_{1}\kappa_{2,j}}\Biggr].$
Theorem 2.2 ensures the rationality of SMC for posterior sampling derived from a Gaussian prior. The proof is provided in 5.4.4. Under an appropriate problem setup, this error bound attains its minimum at some finite $J>1$. Thus, the error bound of the multi-layer SMC is smaller than that of the trivial single-layer case, which illustrates the necessity of developing the SMC algorithm. For more details, see 5.4.4. In the next subsection, we will refine the pCN algorithm, which is also designed for posteriors derived from a Gaussian prior. Without Theorem 2.2, using pCN as a transition kernel would lack theoretical justification.
@@
The key components of this kernel are the proposal $Q(u,dv)$ and the acceptance rate function $a(u,v)$. To sample from the target measure $\rho$, which is absolutely contiounous with respect to prior $\mu_{0}$, we must select $Q(u,dv)$ and $a(u,v)$ such that $P(u,dv)$ defined by (2.17) preserves $\rho$ as follows:
$\int_{u\in\mathcal{H}}P(u,dv)\rho(du)=\rho(dv).$ (2.18)
@@
$a_{\text{\tiny MH}}=\min\left\{1,\frac{\rho(dv)Q(v,du)}{\rho(du)Q(u,dv)}\right\},$ (2.19)
so that $P(u,dv)$ preserves $\rho$, provided that $\rho(du)Q(u,dv)$ and $\rho(dv)Q(v,du)$ are equivalent as measures on $(\mathcal{H}\times\mathcal{H},B(\mathcal{H})\times B(\mathcal{H}))$. We will refer to a Metropolis-Hastings algorithm as well-defined if this condition of absolute continuity is satisfied Cui2016 . Hence, we only need to select the proposal kernel $Q(u,dv)$ and demonstrate the equivalence between $\rho(du)Q(u,dv)$ and $\rho(dv)Q(v,du)$.
@@
where $\mathcal{N}_{0,\mathcal{C}}$ denotes the Gaussian measure $\mathcal{N}(0,\mathcal{C})$. Note that (2.20) is a drifted version Pinski2015 of the pCN proposal. The special case without drift (i.e., when $m=0$) is more frequently discussed Andrew_review , Beskos2015SC . This kernel allows for deriving a simple acceptance rate function
$a_{\text{pCN}}(u,v)=\min\left\{1,\exp(\Phi(u)-\Phi(v))\right\},\quad\frac{d\rho}{d\mu_{0}}(u)\propto e^{-\Phi(u)}.$ (2.21)
As an MCMC sampling method, the irreducibility of the pCN algorithm is of fundamental importance, as it is closely related to ergodicity, the law of large numbers, and central limit theorems. For MCMC methods arising from discretizations of continuous-time dynamical systems, irreducibility may sometimes depend on the specific discretization scheme. A notable example is Hamiltonian Monte Carlo Alain2020 . Due to intrinsic properties of infinite-dimensional Hilbert spaces, the pCN algorithm generally loses irreducibility. Nevertheless, ergodicity, the law of large numbers, and central limit theorems still hold. Since a detailed discussion of these issues is not the main focus of this paper, we refer the reader to 5.1 for more details.
The pCN method will serve as the transition kernel for SMC in numerical experiments, constituting an accurate posterior sampling algorithm for comparison. On the other hand, to prove the convergence of our algorithm (in Section 2.3), we generalize the pCN method to the case with a Gaussian mixture proposal, and prove that it is also well-defined in an infinite-dimensional space.
@@
This proposal has been carefully constructed, and it will yield a well-defined Metropolis-Hastings algorithm at the end of this subsection. We can obtain this proposal from the Crank-Nicolson discretization of a Langevin system, see 5.2. This proposal degenerates to the pCN proposal with a non-zero mean prior Pinski2015 when $M=1$, and to the proposal in the Gaussian mixture independence sampler Feng2018 when $\beta=1$. We will estimate the parameters of a Gaussian mixture from samples by the method from Feng2018 , which is established in function space. And the resulting Gaussian mixture is also equivalent to the prior measure $\mu_{0}$.
@@
$a(u,v)=\min\left\{1,\frac{\rho(dv)Q(v,du)}{\rho(du)Q(u,dv)}\right\}.$ (2.25)
Here the accept rate function $a(u,v)$ is well-defined in infinite-dimensional Hilbert space if and only if the Radon-Nikodym derivative exists, that is, $\rho(dv)Q(v,du)$ is absolutely continuous with respect to $\rho(du)Q(u,dv)$, as measures on $\mathcal{H}\times\mathcal{H}$. We will first illustrate the well-definiteness of the measure $\mu_{0}(dv)Q(v,du)$ and then show their equivalence.
@@
$\mu_{0}(du)Q(u,dv)=$ $\sum_{j=1}^{M}w_{j}\rho_{j}(du,dv),\quad\rho_{j}=\mathcal{N}\left(m_{j},\mathcal{V}_{j}\right),$ (2.26)
$\mu_{0}(dv)Q(v,du)=$ $\sum_{j=1}^{M}w_{j}\rho_{j}^{\prime}(du,dv),\quad\rho_{j}^{\prime}=\mathcal{N}\left(m_{j}^{\prime},\mathcal{V}_{j}^{\prime}\right).$ (2.27)
@@
The proof of theorem 2.3 is given in 5.4.5. Furthermore we can prove the equivalence between $\mu_{0}(dv)Q(v,du)$ and $\mu_{0}(du)Q(u,dv).$
@@
Under the assumptions of theorem 2.3 we have that the operator $\mathcal{V}_{j}^{-1/2}\mathcal{V}_{j}^{\prime}\mathcal{V}_{j}^{-1/2}$-$\mathcal{I}$ is a Hilbert-Schmidt operator, and $\mu_{0}(du)$ $Q(u,dv)$, $\mu_{0}(dv)Q(v,du)$ are equivalent. In fact, all of their components are equivalent. Furthermore, $\rho(du)Q(u,dv)$ and $\rho(dv)Q(v,du)$ are equivalent.
@@
The equivalence between the two measures in this theorem is crucial (see the proof in 5.4.6), directly affecting whether the algorithm possesses dimension independence. A simple example is the comparison between the pCN and the random walk transition kernel:
@@
where just a slight difference in coefficients can lead to significant changes in the algorithm’s acceptance rate as the discrete dimension increases Andrew_review . Moreover, if we require a proposal of the form given in Equation (2.24) without the assumption $\gamma=\sqrt{1-\beta^{2}}$ to yield a well-defined acceptance rate function, then the only option is $\beta^{2}+\gamma^{2}=1$. Discarding the meaningless negative solutions, the generalization we provide is unique. The proof of uniqueness is given in 5.4.7.